OpenStudy (anonymous):

Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right.

OpenStudy (ajprincess):

Do the opposite of the given operations to find the equation. Let x be a variable. Multiply it by 4. u get 4x. subtract 7 from it. u get 4x-7. Equate it to zero to get the eqn. 4x-7=0 add 7 to both sides 4x-7+7=0+7 4x=7 divide both sides by 4 4x/4=7/4 x=7/4